OpenStudy (anonymous):

HELP Freddie is at chess practice waiting on his opponent's next move. He notices that the 4-inch-long minute hand is rotating around the clock and marking off time like degrees on a unit circle. Part 1: How many radians does the minute hand move from 3:35 to 3:55? (Hint: Find the number of degrees per minute first.) Part 2: How far does the tip of the minute hand travel during that time? Part 3: How many radians on the unit circle would the minute hand travel from 0 degrees if it were to move 3pi inches? Part 4: What is the coordinate point associated with this radian measure?

OpenStudy (anonymous):

11 min for the first. the rest I do not know.

OpenStudy (anonymous):

How do you find that? I have to type out how I figured it out. @Delila75729493894

OpenStudy (misty1212):

HI!!! there are 60 minutes in an hour, and 380 degrees in a circle how many degrees per minute?

OpenStudy (anonymous):

@misty1212 would I divide 380/60?

OpenStudy (misty1212):

hint if you need it, "per" means "divided by"

OpenStudy (misty1212):

yes

OpenStudy (anonymous):

6.3

OpenStudy (misty1212):

oops except i made a mistake, it is 360 in a circle,so \)360\div 60=6\)

OpenStudy (misty1212):

\[360\div 60=6\] my fault

OpenStudy (misty1212):

so 6 per minute, five minutes, you have \(6\times 5=30^\circ\) in five minutes

OpenStudy (anonymous):

okay

OpenStudy (misty1212):

now that we have it in degrees we need to convert to radians like it asks

OpenStudy (misty1212):

so \[\frac{30}{180}=\frac{1}{6}\] the radian measure is \(\frac{\pi}{6}\)

OpenStudy (misty1212):

for the next one multiply \(\frac{\pi}{6}\times 4\)

OpenStudy (anonymous):

2.09?

OpenStudy (misty1212):

leave it as \[\frac{2\pi}{3}\]

OpenStudy (misty1212):

for the next one on the unit circle the angle measure is the same as the arc length

OpenStudy (misty1212):

so it would travel \(3\pi\) radians

OpenStudy (anonymous):

I think I'm understanding, its virtual school and I suck at it.

OpenStudy (misty1212):

the coordinates would be \((-1,0)\) gotta run to class

OpenStudy (anonymous):

Thank you though!

OpenStudy (anonymous):

how did you find the coordinate??? @misty1212 please help

OpenStudy (anonymous):

@inspirechey I found the answer. it's (-2.8,-2.8).

OpenStudy (anonymous):

This doesn't seem right..

OpenStudy (anonymous):

Why would you multiply by 5 when it's 20 minutes?